# Differentially Private Matchings

### v2

@@ 2.1 Differential Privacy
We begin with the basic definitions of differential privacy for graphs. Two graphs $G$ and $G^{\prime}$ are said to be edge-neighboring if they differ in one edge. They are said to be node-neighboring if they differ in one node (and all edges incident to that node).
@@
$\Pr[\mathcal{A}(G)\in S]\leq e^{\varepsilon}\Pr[\mathcal{A}(G^{\prime})\in S]+\delta$
@@
Definition 2.2 (Local Randomizer (LR) [DLRSSY22]).
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge DP algorithm that takes as input the set of its neighbors $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$. In other words,
@@
Definition 2.3 (Local Edge Differential Privacy (LEDP) [DLRSSY22]).
@@
Given an input graph $G=(V,E)$, algorithms in the billboard model produces a public billboard. Then, each node $v\in V$ in the graph deduces the portion of the explicit solution that $v$ participates in.
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then the explicit solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
@@
We now define the implicit solutions that our algorithms will post to the billboard. Note that by the definition of $\varepsilon$-LEDP, the entire transcript is public and so without loss of generality is posted to the billboard. But the particular information that each node will use to produce its explicit solution is the following.
@@
Given a graph $G=(V,E)$, an implicit solution is a collection $\mathcal{S}=\{S_{v}\}_{v\in V}$ where each $S_{v}\subseteq V$. An implicit solution $\mathcal{S}$ defines an implicit graph $H(\mathcal{S})=(V,E^{\prime})$ where $E^{\prime}=\{\{u,v\}:u\in S_{v}\lor v\in S_{u}\}$. The degree of an implicit solution $\mathcal{S}$ is the maximum degree in the graph $(V,E\cap E(H(\mathcal{S})))$, i.e., the maximum degree in the graph which is the intersection of $G$ and $H(\mathcal{S})$.
In other words, for each vertex $v$ we have a subset of nodes $S_{v}$. Think of $S_{v}$ as “potential matches” for $v$. Then $H(\mathcal{S})$ is the graph obtained by adding an edge if either endpoint contains the other as a potential match. We get a third graph by intersecting the implicit graph with the true graph, and the maximum degree in this graph is what we call the degree of the solution. From a billboard/LEDP perspective, this means that any node $v$ can locally decode an explicit solution from the implicit solution since it can perform the intersection of $H(\mathcal{S})$ with its own neighborhood.
@@
Given a complete graph $G=(V,E)$, each edge in the graph is given a binary weight of 00 or $1$. Edge-neighbors are two graphs $G$ and $G^{\prime}$ where the weight of exactly one edge differs. Node-neighbors are two graphs where the weights of all edges adjacent to exactly one node differ.
@@ 2.2 Differential Privacy Tools
1 Input: Graph $G$, adaptive queries $\{\vec{f}_{1},\dots,\vec{f}_{n}\}$, threshold vector $\vec{T}$, privacy $\varepsilon$, $\ell_{1}$-sensitivity $\Delta$.
@@
Lemma 2.17 (Privacy Amplification via Subsampling).
@@
3 Lower Bound for Explicit Matchings
@@
We now prove Theorem 1.1 which is restated below for convenience: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 2.7), the required error is essentially linear, even if we allow for the algorithm to output a $b$-matching for very large $b$.
See 1.1
Given a graph $G=(V,E)$, let $\nu(G)$ denote the size of a maximal matching in $G$ (while this is not fully well-defined, all graphs on which we use this definition will have the property that all maximal matchings have the same size, and so are actually maximum matchings). Given a (randomized) algorithm $\mathcal{A}$ which outputs some graph $H=(V,E_{H})$ when run on input graph $G=(V,E)$, let $\nu_{\mathcal{A}}(G)=\nu((V,E\cap E_{H}))$ denote the size of a maximal matching in the graph consisting of edges that are in both $G$ and $H$.
@@
At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that the marginal probability of choosing each non-edge is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
3.1 Proof of Theorem 1.1
@@
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a $b$-matching on $V$ (i.e. a graph with vertex set $V$ in which every vertex has degree at most $b$). For $G\in\mathcal{M}$ and $\{u,v\}\in\binom{V}{2}$, let $p(\mathcal{A},G,\{u,v\})$ denote the probability that $\mathcal{A}(G)$ contains $\{u,v\}$. Let $\nu_{\mathcal{A}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G)]$ denote the worst-case expected utility of $\mathcal{A}$ on $\mathcal{M}$. Note that for $G\in\mathcal{M}$, every subset of $E(G)$ is a matching. Hence $\text{E}[\nu_{\mathcal{A}}(G)]=\text{E}[|E(G)\cap\mathcal{A}(G)|]\geq\nu_{\mathcal{A}}$ for all $G\in\mathcal{M}$. By linearity of expectations, this is equivalent to saying that $\sum_{\{u,v\}\in E(G)}p(\mathcal{A},G,\{u,v\})\geq\nu_{\mathcal{A}}$ for all $G\in\mathcal{M}$.
@@
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}$ on $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$, and moreover satisfies the following symmetry property: $p(\mathcal{A}^{\prime},G,\{u,v\})=p(\mathcal{A}^{\prime},G,\{u^{\prime},v^{\prime}\})$ for all $\{u,v\},\{u^{\prime},v^{\prime}\}\not\in E(G)$.
@@
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $\pi(H)$ denote the resulting $b$-matching. We then “undo” $\pi$ to get the $b$-matching $H=\{\{u,v\}:\{\pi(u),\pi(v)\}\in\pi(H)\}$. We return $H$.
@@
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This is straightforward, since it simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. Slightly more formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
Now we claim that $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Let $\alpha_{\mathcal{A}}(G)$ denote the expected utility of $\mathcal{A}$ on $G\in\mathcal{M}$, and similarly let $\alpha_{\mathcal{A}^{\prime}}(G)$ denote the expected utility of $\mathcal{A}^{\prime}$ on $G$. Note that by construction, there is a bijection between $\pi(H)\cap E(\pi(G))$ and $H\cap E(G)$, and hence $\nu_{\mathcal{A}^{\prime}}(G)=\nu_{\mathcal{A}}(\pi(G))$. Since $\pi$ was chosen uniformly at random, we know that $\pi(G)$ is distributed uniformly among $\mathcal{M}$. Hence for any $G\in\mathcal{M}$ we have that
$\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]$ $=\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}}(G^{\prime})]\geq\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\nu_{\mathcal{A}}=\nu_{\mathcal{A}}.$
@@
Sicne this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}=\min_{G\in\mathcal{M}}\text{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
@@
We now prove the symmetry property. Note that there are $n!$ permutations in $S_{V}$, but only $\frac{n!}{2^{n/2}(n/2)!}$ different perfect matchings (i.e., $|\mathcal{M}|=\frac{n!}{2^{n/2}(n/2)!}$). This is because $\pi(G)$ can equal $\pi^{\prime}(G)$ for many for many $\pi\neq\pi^{\prime}$. In particular, it is not hard to see that |$|\{\pi\in S_{V}:\pi(G)=G^{\prime}\}|=2^{n/2}(n/2)!$ for all $G,G^{\prime}\in\mathcal{M}$. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. It is not hard to see that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$p(\mathcal{A}^{\prime},G,\{u,v\})$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}p(\mathcal{A},G^{\prime},\{\pi(u),\pi(v)\})$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}p(\mathcal{A},G^{\prime},\{\pi(u),\pi(v)\})$
$=\frac{1}{|\mathcal{M}|}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})}p(\mathcal{A},G^{\prime},\{u^{\prime},v^{\prime}\}).$
@@
Hence $p(\mathcal{A}^{\prime},G,\{u,v\})$ is actually independent of $\{u,v\}$, and thus for all $\{u,v\},\{u^{\prime},v^{\prime}\}\not\in E(G)$ we have that $p(\mathcal{A}^{\prime},G,\{u,v\})=p(\mathcal{A}^{\prime},G,\{u^{\prime},v^{\prime}\})$ as claimed. ∎
@@
Lemma 3.1 implies that if we can prove an upper bound on the expected utility) for DP algorithms which obey the symmetry property, then the same upper bound applies to all DP algorithms. We we will now prove such a bound.
@@
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a $b$-matching, and satisfies the symmetry property of Lemma 3.1. Then
$\nu_{\mathcal{A}}\leq e^{2\varepsilon}b+\delta n.$

### v3

@@ 3.1 Differential Privacy
We begin with the basic definitions of differential privacy for graphs. Two graphs $G$ and $G^{\prime}$ are said to be edge-neighboring if they differ in one edge, i.e., there are vertices $u,v\in V(G)=V(G^{\prime})$ such that $\{uv\}=(E(G)\setminus E(G^{\prime}))\cup(E(G^{\prime})\setminus E(G))$. They are said to be node-neighboring if they differ in all edges incident to a particular node. Differentially private algorithms on edge-neighboring and node-neighboring graphs satisfy edge-privacy [NRS07] and node-privacy [BBS13, CZ13, kasiviswanathan2013analyzing], respectively.
@@
$\Pr[\mathcal{A}(G)\in S]\leq e^{\varepsilon}\cdot\Pr[\mathcal{A}(G^{\prime})\in S]+\delta$
@@
Definition 3.2 (Local Randomizer (LR) [DLRSSY22, kasiviswanathan2011can]).
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge DP algorithm that takes as input the set of its neighbors, $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$. In other words,
@@
Definition 3.3 (Local Edge Differential Privacy (LEDP) [DLRSSY22, JMN019]).
@@
Given an input graph $G=(V,E)$, algorithms in the billboard model produces a public billboard $B$. Then, each node $v\in V$ in the graph deduces the portion of the explicit solution that $v$ participates in as a function of $B$ and its private neighborhood $N(v)$.
In particular, one can easily show that if every node processes the information contained in the public billboard using a deterministic algorithm, then explicitly revealing the solution contained at every node will be $\varepsilon$-edge DP (with respect to edge-neighboring graphs) except for the two nodes adjacent to the edge that differs between neighboring graphs. Morally speaking, such a set of explicit solutions does not leak any additional private information since the nodes that are endpoints to the edge that differs already know that this edge exists. Hence, we decouple the privacy definition from the solution release model and say an algorithm is $\varepsilon$-LEDP in the billboard model if the public billboard is produced via a $\varepsilon$-LEDP algorithm and the explicit solutions obtained by every node is via a deterministic algorithm at each node.
@@
We now define the implicit solutions that our algorithms will post to the billboard. Note that by the definition of $\varepsilon$-LEDP, the entire transcript is public and so, without loss of generality, is posted to the billboard. But the particular information that each node will use to produce its explicit solution is the following.
@@
Given a graph $G=(V,E)$, an implicit solution is a collection $\mathcal{S}=\{S_{v}\}_{v\in V}$ where each $S_{v}\subseteq V$. An implicit solution $\mathcal{S}$ defines a graph $H(\mathcal{S})=(V,E^{\prime})$ where $E^{\prime}=\{\{u,v\}:u\in S_{v}\lor v\in S_{u}\}$. The degree of an implicit solution $\mathcal{S}$ is the maximum degree in the implicit graph $(V,E\cap E(H(\mathcal{S})))$, i.e., the maximum degree in the graph which is the intersection of $G$ and $H(\mathcal{S})$.
In other words, for each vertex $v$ we have a subset of nodes $S_{v}$. Think of $S_{v}$ as “potential matches” for $v$. Then, $H(\mathcal{S})$ is the graph obtained by adding an edge if either endpoint contains the other as a potential match. We obtain a third graph by intersecting the implicit graph with the true graph, and the maximum degree in this graph is the degree of the solution. From a billboard/LEDP perspective, this means that any node $v$ can locally decode an explicit solution from the implicit solution since it can perform the intersection of $H(\mathcal{S})$ with its own neighborhood.
@@
Given a complete graph $G=(V,E)$, each edge in the graph is given a binary weight of 0 or $1$. Edge-neighbors are two graphs $G$ and $G^{\prime}$ where the weight of exactly one edge differs. Node-neighbors are two graphs where the weights of all edges adjacent to exactly one node differ.
Remark 3.8.
For a given implicit solution $\mathcal{S}$ (Definition 3.6), we remark that the graph $H(\mathcal{S})$ is a valid explicit solution in the 0/1-weight model (Definition 3.7). We define $\mathcal{S}$ as a collection of local neighborhoods to highlight the fact that our algorithms are local.
Suppose that $\mathcal{S}$ is the output of an edge-DP approximate maximum matching algorithm, then $H(\mathcal{S})$ would be a solution that attains the same approximation ratio in the 0/1-weight model. On one hand, our implicit matching algorithms (Theorem 2.2) ensure that the implicit graph $H(\mathcal{S})\cap G$ has $O(\log n)$ maximum degree. On the other hand, our explicit solution lower bound (Theorem 2.1) ensures that $H(\mathcal{S})$ must have a large $\Omega(n)$ maximum degree. This demonstrates that implicit solutions are necessary to obtain meaningful algorithmic solutions.
@@ 3.2 Differential Privacy Tools
1Input: Graph $G$, adaptive queries $\{\vec{f}_{1},\dots,\vec{f}_{n}\}$, threshold vector $\vec{T}$, privacy parameter $\varepsilon$, $\ell_{1}$-sensitivity $\Delta$.
@@
Lemma 3.18 (Privacy Amplification via Subsampling [li2012sampling]).
@@
4 Lower Bounds for Explicit Solutions via Symmetrization
In this section, we introduce our novel symmetry argument technique for proving lower bounds for graph differential privacy. As previously mentioned, this is inspired by arguments from distributed computation. Connections between differential privacy and other areas have previously been studied [bun2023stability, dwork2015adaptive, beimel2022adaptive, feldman2020memorization]. We hope that our work can lead to more connections between DP and distributed algorithms, a direction so far underexplored.
@@ 4 Lower Bounds for Explicit Solutions via Symmetrization
Our goal is to prove Theorem 2.1 which shows, intuitively, that any explicit solution (when provided $0/1$ weights for the graph) for the maximum matching problem must have large error. More specifically, Theorem 2.1 as restated below states: in the basic 0/1 weight $(\varepsilon,\delta)$-edge DP model (Definition 3.7), the required error is essentially linear, even if we allow for the algorithm to output a $b$-matching for very large $b$.
See 2.1
We will prove a more general lower bound for synthetic graph generation (Definition 4.1), which can be interpreted as a relaxation of the explicit matchings. As we will see, this is a stronger result than Theorem 2.1 since the synthetic graph problem has less stringent constraints than matchings. In Section 4.1, we formally define this problem and prove its hardness. Then, we prove Theorem 2.1 as a straightforward corollary in Section 4.2.
4.1 Lower Bound on DP Synthetic Graph Generation
We now formalize our symmetry argument and use it to prove the following lower bound (Theorem 4.3), which is a relaxation of our explicit matching problem. We will use Theorem 4.3 to prove Theorem 2.1 in Section 4.2. Note that our symmetry argument is also used to prove the lower bound presented in Theorem 2.5.
Definition 4.1 (Synthetic Graph Generation).
Let $G=(V,E)$ be an input graph and $b\in[0,n]^{V}$ be vertex degree bounds. The corresponding synthetic graph problem is given by the following stochastic program. This program states: given (potentially randomized) variables $x_{uv}$, one for each pair of vertices $u\neq v\in V$, we want to maximize the expected sum of all variables corresponding to edges in $E$, subject to the constraints that the expected sum of all variables associated with any vertex $v$ is at most $b_{v}$.
$\max$ $\sum_{uv\in E(G)}\mathbb{E}[x_{uv}]$
$\sum_{u\neq v}\mathbb{E}[x_{uv}]\leq b_{v}$ $\forall v\in V$
$\mathbf{x}$ $\in[0,1]^{\binom{V}{2}}$
Synthetic graph generation (Definition 4.1) is interesting due to the fact that many LEDP mechanisms, such as randomized response, produce dense graph outputs. Such dense graph outputs are undesirable both due to the large error as well as slow processing times. Our lower bound below essentially states that such dense outputs are unavoidable if one wants to preserve differential privacy. To the best of our knowledge, while sparsity lower bounds for specific algorithm such as randomized response have been studied [epasto2024smooth] under pure $\varepsilon$-DP, Theorem 4.3 is the first algorithm-agnostic lower bound for synthetic graph generation under approximate $(\varepsilon,\delta)$-DP.
Remark 4.2.
In the case that $\mathbf{x}\in\{0,1\}^{\binom{V}{2}}$ is required to be integral, we can interpret Definition 4.1 as outputting a graph $H$ maximizing $\mathbb{E}[|E(H)\cap E(G)|]$ subject to degree constraints in expectation $\mathbb{E}[\deg_{H}(v)]\leq b_{v}$. This problem essentially asks for a synthetic graph with maximum overlap with a given input graph $G$ and has degree constraints on all vertices.
This can be viewed as a relaxation of the 0/1-weight matching problem in the case $E(G)$ is a matching (e.g., the maximum degree of $G$ is 1), in the sense that the output of any (randomized) 0/1-weight matching algorithm on such a graph $G$ is a feasible solution to the synthetic graph problem with input $G$.
Having defined the synthetic graph problem, we are ready to state the lower bound Theorem 4.3 below. In the below theorem, a solution $S\in[0,1]^{{V\choose 2}}$ is a $(\gamma,\beta)$-approximation of the solution to the Synthetic Graph Problem (Definition 4.1) if $\sum_{uv\in E}\mathbb{E}[S_{uv}]\geq\frac{|E|}{\gamma}-\beta$.
Theorem 4.3 (Lower Bound on Private Synthetic Graph Approximation).
Let $\mathcal{A}$ be an algorithm which satisfies $(\varepsilon,\delta)$-edge DP and always outputs a solution $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ to the synthetic graph problem (Definition 4.1) parameterized by graph $G$ and degree bounds $b$ with average degree bound $\bar{b}\coloneqq\frac{1}{n}\sum_{v\in V}b_{v}$. If $\mathcal{A}$ outputs a $(\gamma,\beta)$-approximation (even in expectation), then $\gamma,\beta$ must satisfy
$e^{4\varepsilon}\bar{b}+2\delta n\geq\frac{n}{2\gamma}-\beta\,.$
4.1.1 Proof of Theorem 4.3
@@ 4.1 Lower Bound on DP Synthetic Graph Generation
We now present our proof of Theorem 4.3. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument, we will instead use a novel symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that $\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ for each non-edge $uv\notin E$ is identical. This makes it easy to argue that high-utility algorithms cannot satisfy differential privacy.
@@
We begin with a description of our inputs and some more notation. Given $n$ nodes $V$ for an even integer $n>0$, let $\mathcal{M}$ denote the set of all perfect matchings on $V$. Let $\mathcal{A}$ be an $(\varepsilon,\delta)$-differentially private algorithm which always returns a fractional solution $X_{G}^{\mathcal{A}}\in[0,1]^{\binom{V}{2}}$ as defined above (Definition 4.1). For $G\in\mathcal{M}$ and $uv\in\binom{V}{2}$, let $\mathbb{E}[X_{G,uv}^{\mathcal{A}}]$ denote the expected fractional value that $\mathcal{A}(G)$ assigns to $uv$.
@@
If $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-DP and has expected utility at least $\nu_{\mathcal{A}}$ over $\mathcal{M}$, then there is an algorithm $\mathcal{A}^{\prime}$ which also satisfies $(\varepsilon,\delta)$-DP and has $\nu_{\mathcal{A}^{\prime}}\geq\nu_{\mathcal{A}}$. Moreover $\mathcal{A}^{\prime}$ satisfies the following symmetry property: $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\mathbb{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ for all $uv,ab\not\in E(G)$ when executed on input $G\in\mathcal{M}$.
@@
Consider the following algorithm $\mathcal{A}^{\prime}$. We first choose a permutation $\pi$ of $V$ uniformly at random from the space $S_{V}$ of all permutations of $V$. Let $\pi(G)$ denote the graph isomorphic to $G$ obtained by applying this permutation, i.e., by creating an edge $\{\pi(u),\pi(v)\}$ for each edge $\{u,v\}\in E(G)$. We run $\mathcal{A}$ on $\pi(G)$, and let $X_{\pi(G)}^{\mathcal{A}}$ denote the resulting fractional solution for $\pi(G)$. We then “undo” $\pi$ to get the fractional solution $X_{G,uv}^{\mathcal{A}^{\prime}}=X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}$. We return $X_{G}^{\mathcal{A}^{\prime}}$.
@@
We first claim that $\mathcal{A}^{\prime}$ is $(\varepsilon,\delta)$-DP. This can be seen since $\mathcal{A}^{\prime}$ simply runs $\mathcal{A}$ on a uniformly random permuted version of $G$, and then does some post-processing. More formally, for any neighboring graphs $G$ and $G^{\prime}$ and for any $\pi\in S_{V}$, the graphs $\pi(G)$ and $\pi(G^{\prime})$ are also neighboring graphs. Hence the differential privacy guarantee of $\mathcal{A}$ implies that running $\mathcal{A}$ on $\pi(G)$ and $\pi(G^{\prime})$ is $(\varepsilon,\delta)$-DP, and then “undoing” $\pi$ is simply post-processing. So $\mathcal{A}^{\prime}$ has the same privacy guarantees as $\mathcal{A}$.
$\mathbb{E}[\nu_{\mathcal{A}^{\prime}}(G)]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\mathbb{E}[\nu_{\mathcal{A}}(\pi(G))]\geq\frac{1}{n!}\sum_{\pi\in S_{V}}\nu_{\mathcal{A}}=\nu_{\mathcal{A}}\,.$
@@
Since this is true for all $G\in\mathcal{M}$, we have that $\nu_{\mathcal{A}^{\prime}}\coloneqq\min_{G\in\mathcal{M}}\mathbb{E}[\nu_{\mathcal{A}^{\prime}}(G)]\geq\nu_{\mathcal{A}}$, as claimed.
@@
We now prove the key symmetry property. Fix $G,G^{\prime}\in\mathcal{M}$, and let $\Pi(G,G^{\prime})=\{\pi\in S_{V}:\pi(G)=G^{\prime}\}$. Note that if we draw a permutation $\pi$ uniformly at random from $\Pi(G,G^{\prime})$, then for any $\{u,v\}\not\in E(G)$, the pair $\{\pi(u),\pi(v)\}$ is uniformly distributed among the set $\{\{u^{\prime},v^{\prime}\}\not\in E(G^{\prime})\}$. Thus we have that
$\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ $=\frac{1}{n!}\sum_{\pi\in S_{V}}\mathbb{E}[X_{\pi(G),\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{\pi\in\Pi(G,G^{\prime})}\mathbb{E}[X_{G^{\prime},\pi(u)\pi(v)}^{\mathcal{A}}]$
$=\frac{1}{n!}\sum_{G^{\prime}\in\mathcal{M}}\sum_{ab\not\in E(G^{\prime})}\mathbb{E}[X_{G^{\prime},ab}^{\mathcal{A}}].$
@@
In particular, $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]$ is actually independent of $u,v$, and thus for all $uv,ab\not\in E(G)$, we have that $\mathbb{E}[X_{G,uv}^{\mathcal{A}^{\prime}}]=\mathbb{E}[X_{G,ab}^{\mathcal{A}^{\prime}}]$ as claimed. ∎
@@
Lemma 4.4 implies that if we can prove an upper bound on the expected utility for DP algorithms that obey the symmetry property, then the same upper bound applies to all DP algorithms. We will now prove such a bound.
@@
Let $\mathcal{A}$ be an algorithm which is $(\varepsilon,\delta)$-DP, always outputs a synthetic graph solution (Definition 4.1), and satisfies the symmetry property of Lemma 4.4. Then, for $\bar{b}\coloneqq\frac{1}{n}\sum_{v\in V}b_{v}$ and $n\geq 4$,
$\nu_{\mathcal{A}}\leq e^{4\varepsilon}\bar{b}+\delta n\,.$